Circular No. 6254

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995ak IN IC 1844 C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery,

on a Tech Pan film taken with the OCA Schmidt telescope on Oct. 27.06 UT, of a supernova (V ∼ 16.4) at α = 2^{h}45^{m}48^{s}.83, δ = +3^{o}13′50′′.1 (equinox 2000.0, which is 8′′.5 west and 1′′.0 north of the center of IC 1844 (and situated to the south of a large, diffuse and complex blue area of the host galaxy). A nearby star (V ∼ 16.0) has position end figures 53^{s}.34, 14′36′′.2. SN 1995ak does not appear on the original Palomar Sky Survey prints, and no star is seen at this position on an OCA film obtained in 1992.

E. Cappellaro and M. Turatto, Osservatorio di Asiago, report their observations of SN 1995ak obtained on Oct. 31.0 UT with the 1.8-m telescope (+ Boller & Chivens spectrograph + CCD): “A preliminary reduction of the spectrum shows prominent Si II and S II lines measured at 628.5, 578.9, 559.5, and 542.4 nm on a relatively blue continuum. Once the spectrum is corrected for the redshift of the parent galaxy (z = 0.023), as deduced from the strong narrow Hα + [N II] emission line superimposed on the supernova spectrum, we note a close resemblance to spectra of SN 1994D at maximum (Patat et al., MNRAS, in press), confirming that the candidate is a type-Ia supernova at maximum. The expansion velocity of the photosphere measured from the minimum of the Si II absorption is 9980 km/s.”

NGC 3786 B. Nelson, Department of Astronomy, University of California at Los

Angeles (UCLA), writes: “NGC 3786 (= Mrk 744 = Arp 294 = UGC 6621 = VV 228b), a relatively unknown Seyfert galaxy, has recently undergone a major outburst at both optical and infrared wavelengths, as observed with the UCLA 0.61-m telescope. Beginning roughly in 1994 Dec., the optical outburst lasted until 1995 June, with a peak V flux in April that was 30 percent brighter than pre- and post-outburst observations. In the infrared at K′ (2.1 μm), a similar, longer-lived outburst was observed, with a 40percent increase in brightness, and a peak occurring roughly a month after that at V. We are interested in finding other observations of this galaxy at any wavelength, especially covering the time period from 1994 to date. Please direct all correspondence to nelson@eggneb.astro.ucla.edu.”

1995 October 31 Daniel W. E. Green